# Black-Hole Cartography

### v1

@@ Abstract
Quasinormal modes (QNMs) are usually characterized by their time dependence; oscillations at specific frequencies predicted by black hole (BH) perturbation theory. QNMs are routinely identified in the ringdown of numerical relativity waveforms, are widely used in waveform modeling, and underpin key tests of general relativity and of the nature of compact objects; a program sometimes called BH spectroscopy. Perturbation theory also predicts a specific spatial shape for each QNM perturbation. For the Kerr metric, these are the ($s=-2$) spheroidal harmonics. Spatial information can be extracted from numerical relativity by fitting a feature with known time dependence to all of the spherical harmonic modes, allowing the shape of the feature to be reconstructed; a program initiated here and that we call BH cartography. Accurate spatial reconstruction requires fitting to many spherical harmonics and is demonstrated using highly accurate Cauchy-characteristic numerical relativity waveforms. The loudest QNMs are mapped, and their reconstructed shapes are found to match the spheroidal harmonic predictions. The cartographic procedure is also applied to the quadratic QNMs – nonlinear features in the ringdown – and their reconstructed shapes are compared with predictions from second-order perturbation theory. BH cartography allows us to determine the viewing angles that maximize the amplitude of the quadratic QNMs, an important guide for future searches, and is expected to lead to an improved understanding of nonlinearities in BH ringdown.
@@ I Introduction
The ground-based gravitational wave (GW) observatories LIGO [1] and Virgo [2] have already observed the merger of dozens of high-mass binary black holes (BHs) [3] and the current observing run is expected to yield many more observations [4, 5]. The final stage of the merger process, known as the ringdown, is associated with the remnant object settling down into its final state. The ringdown can be described by perturbation theory (PT) which predicts the existence of damped sinusoidal oscillations, known as quasinormal modes (QNMs) [6].
QNMs can be used directly for parameter estimation, especially for high-mass systems where only the late inspiral is observed [7, 8], or indirectly when incorporated as part of inspiral-merger-ringdown waveform models [9, 10, 11]. However, the most important application of QNMs is to tests of general relativity (GR), fundamental physics, and the nature of extremely compact objects. QNMs have been used to test Hawking’s area theorem and the consistency of the inspiral and ringdown signals [12, 13]. The no-hair theorem states that BHs should be completely characterized by their mass and spin (charge is excluded for astrophysical BHs). When multiple QNMs are detected, consistency between their frequencies allows us to test this in a program called BH spectroscopy [14, 15, 16]. This requires the clear identification of at least two QNMs and this has been the subject of much discussion in the literature [17, 18, 19, 20, 21]. However, to date, all observations are consistent with the standard picture of a vacuum Kerr BH in GR [22].
@@
Figure 1: Left: From NR, the spatial reconstruction of the fundamental QNM ($\ell=m=2$, $n=0$, prograde), normalized to agree with the corresponding spheroidal harmonic, plotted on the sphere in Mollweide projection. This mapping was performed for the 0001 CCE waveform with a ringdown starting $t_{0}=8M$ after the peak of the $\ell=m=2$ mode strain (see Table 1). Right: From PT, the predicted spheroidal harmonic shape. In both panels, the real parts of the functions are plotted and they match well up to a rotation around the $z$-axis. The spatial mismatch between these two functions is $\mathcal{M}_{\alpha}=1.16\times 10^{-9}$ (see Sec. II.3). If the reconstruction is performed at different $t_{0}$, then the reconstruction is observed to rotate around the sphere with angular frequency equal to the real part of the QNM frequency. Because 0001 is non-precessing, there is a symmetry between the ±mplus-or-$\pm m$ modes; the $(\ell,m,n)=(2,-2,0)$ prograde QNM has a similar structure concentrated in the southern hemisphere.
@@
Another motivation for studying QQNMs is their possible relationship to the GW memory effect. Sometimes called non-linear displacement memory, this is the permanent displacement in test particle separation after the passage of a transient GW signal [33, 34, 35, 36, 37]. GW memory has not yet been observed, but is expected to be prominent in the azimuthally-symmetric $m=0$ modes of the signal [38, 39]. Pairs of prograde modes with ±mplus-or-$\pm m$ indices have frequencies that are symmetrically placed in the complex plane and hence source a QQNM with a purely imaginary frequency. These non-oscillatory modes resemble a low-frequency, persistent displacement. Moreover, the combination of parent modes with ±mplus-or-$\pm m$ indices gives rise to QQNMs which are azimuthally symmetric with $m=0$. Together with the inherent nonlinear nature of QQNMs, this hints at a connection which might allow the GW memory effect to be studied using QNMs.
@@
To date, most fits (using either QNMs or QQNMs) have been performed to individual spherical harmonics of the waveform. Alternatively, the GW signal from a particular viewing direction can be fit. These approaches do not use all of the information in the NR simulation. Recently multi-mode, or angle-averaged, fits have been used, fitting a QNM model to all spherical modes simultaneously [40]. By using all the modes (effectively using the signal at all viewing directions) this exploits all of the information available in the NR simulation. Angle-averaged fits required another ingredient from PT: the predicted distribution of the GW radiation in different directions from the source. This distribution will be called the ‘spatial shape’ of the QNM and is a function on the 2-sphere. For the Kerr metric, the QNM shapes are given by the spin-weighted spheroidal harmonics [41]. This information is usually provided in the mode-mixing coefficients that relate the spheroidal and spherical harmonics.
@@
It is also possible to extract the QNM shapes directly from the NR data, without knowledge of the spheroidal harmonics. By fitting a feature with common time dependence to all spherical modes and treating the mode mixing coefficients as a free parameters, the shape can be reconstructed numerically. In analogy to the identification of modes using their time dependence in BH spectroscopy, this approach is called BH cartography.
@@
This process can produce a spatial reconstruction of the shape of any QNM. An example of such a reconstruction is shown in Fig. 1 for the loudest fundamental QNM. Because this reconstruction is performed using the GW signal, it is formally defined on a sphere at future null infinity. However, it reveals the structure of the perturbations that source the GW signal and are located deep in the strong gravitational field, near the photon sphere of the remnant BH.
@@
BH cartography is a novel method for investigating the ringdown. Applied to the linear QNMs, it merely allows us to verify a well-understood prediction of first-order PT, namely the spheroidal harmonics. However, second-order perturbation theory does not make such a clear, unique prediction for the QQNM shapes, leading to different predictions being used in the literature [30, 42]. Reconstructing the shapes of the QQNMs numerically allows us to test the different predictions. These reconstructions also allow us to determine the viewing directions that maximize the amplitude of the QQNMs with implications for their observability. Hence, knowing the quadratic mixing will be important for future waveform modelling.
@@
The layout of this paper is as follows. Sec. II describes the methods used, including a brief introduction to the key elements from BH PT (Sec. II.1), the least-squares fitting procedures to the NR data (Sec. II.2), the new techniques for spatially mapping components of the ringdown signal (Sec. II.3), and the CCE waveforms used in this study (Sec. II.4). In Sec. III the mapping methods are demonstrated by applying them to the linear QNMs which are already well understood from linear PT. This allows us to validate the BH cartography method. Then, Sec. IV applies the mapping to the QQNMs, revealing the spatial structure of the nonlinear perturbations. Concluding remarks are presented in Sec. V. Throughout, natural units are used in which $G=c=1$.
@@ II.1 Ingredients from perturbation theory
GW signals contain two polarizations; these can be conveniently combined into a single complex quantity called simply the GW strain, $h=h_{+}-ih_{\times}$. The radially-scaled GW strain can be expanded in spin-weight $s=-2$ spherical harmonics, ${}_{-2}Y^{\ell m}$,
@@
$rh^{(1)}(t,\theta,\phi)\!=$ $M\!\sum_{\ell,m}\sum_{n=0}^{\infty}\Big{[}C_{\ell mn}e^{-i\omega_{\ell mn}(t-t_{0})}{}_{-2}S_{\ell mn}(\theta,\phi)$
$+C^{\prime}_{\ell mn}e^{-i\omega^{\prime}_{\ell mn}(t-t_{0})}{}_{-2}S_{\ell mn}^{\prime}(\theta,\phi)\Big{]},$ (2)
where it has been assumed, for simplicity, that the frame $(\theta,\phi)$ is adapted to the remnant BH. That is, the spin of the remnant BH points along the $z$-axis of the chosen coordinate system. It is always possible to perform a rotation into such a frame. The superscript (1) in the notation is intended as a reminder that this part of the model for $h$ comes from linear PT. The ringdown is assumed to be the perturbation of a Kerr BH, which is naturally described using the spheroidal harmonics ${}_{-2}S_{\ell mn}$.
Each term in the expansion in Eq. II.1 is a QNM. The indices $\ell\geq 2$ and $|m|\leq\ell$ label angular modes while the index $n\geq 0$ labels the radial-like overtones. For each triple $(\ell,m,n)$ there exists a pair of QNMs with different signs of the real part of their QNM frequency, $\omega^{\prime}_{\ell mn}=-\omega^{*}_{\ell-mn}$. The modes in this pair are sometimes referred to as the regular mode and mirror mode (denoted with a prime). The mirror modes are associated with the modified spheroidal harmonics, ${}_{-2}S^{\prime}_{\ell mn}(\theta,\phi)={}_{-2}S^{*}_{\ell-mn}(\pi-\theta,\phi)$. In Eq. II.1, this pair of regular and mirror modes has been written explicitly however, it will be convenient to distinguish them with another index $p\in\{+,-\}$ for the regular and mirror mode respectively. Therefore, a QNM is uniquely identified by the tuple of four indices $(\ell,m,n,p)$.
The distinction between regular and mirror modes is related to the direction in which the perturbation propagates around the Kerr BH. A mode is said to be prograde (i.e. co-rotating with the BH) if the signs of $m$ and $p$ are the same (if $mp>0$). If $mp<0$, then the mode is said to be retrograde (i.e. counter-rotating with the BH). Modes with $m=0$, along with all QNMs on Schwarzschild BHs, do not fit either the prograde or retrograde label. Typically, in a quasi-circular merger, the prograde modes are excited with significantly larger amplitudes than the retrograde modes.
The QNM model in Eq. II.1 involves many indices. In a moment, when QQNMs are introduced, the number of indices will increase further. Therefore, it is helpful to introduce the following abbreviated notation for the QNM model in Eq. II.1;
@@
The spheroidal harmonics are the natural basis for describing perturbations of Kerr BHs. The spheroidal basis does not align perfectly with the spherical basis. This leads to a phenomenon called mode mixing. The mode mixing coefficients are functions of the BH spin and are defined as
@@
Going beyond linear PT introduces additional terms into the model in Eq. 3. Notable among these are the quadratic QNMs (QQNMs). This second order part of the model, denoted with a superscript (2), is given by
@@ II.2 Least-squares fitting
The free parameters of the QNM model are the QNM complex amplitudes $C_{\alpha}$. The spheroidal harmonics and the QNM frequencies are functions of the final spin magnitude of the remnant BH (the frequencies are also scaled by the remnant mass, $M_{f}$). These were determined from the asymptotic bondi data objects obtained using the scri package [44, 45, 46, 47] with the relevant files from the CCE data, and are given in Table 1.
To perform model fits to data, we use the least-squares fit implemented in the qnmfits package [48]. More details on the single and multi-mode fitting procedure are given in Appendix A. For all simulations, the time $t=0$ is set to be the peak of the amplitude of the 22 mode, $|\mathfrak{h}^{22}(t)|$.
For the NR data used in the fits, if the sums are truncated such that only modes with $\ell\leq\ell_{\rm max}$ are included, the resultant number of spherical harmonics ${}_{-2}Y^{\ell m}$ is $(\ell_{\rm max}+3)(\ell_{\rm max}-1)$. For the QNM models, if the sums are truncated such that all modes with $\ell\leq\ell_{\rm max}$ and $n\leq n_{\rm max}$ are included, the number of QNMs (including both $p=\pm 1$) is $2(\ell_{\rm max}+3)(\ell_{\rm max}-1)(n_{\rm max}+1)$. Unless otherwise stated, all of the results in this paper were obtained using $\ell_{\rm max}=8$, $n_{\rm max}=7$.

### v2

@@ Abstract
Quasinormal modes (QNMs) are usually characterized by their time dependence; oscillations at specific frequencies predicted by black hole (BH) perturbation theory. QNMs are routinely identified in the ringdown of numerical relativity waveforms, are widely used in waveform modeling, and underpin key tests of general relativity and of the nature of compact objects; a program sometimes called BH spectroscopy. Perturbation theory also predicts a specific spatial shape for each QNM perturbation. For the Kerr metric, these are the ($s=-2$) spheroidal harmonics. Spatial information can be extracted from numerical relativity by fitting a feature with known time dependence to all of the spherical harmonic modes, allowing the shape of the feature to be reconstructed; a program initiated here and that we call BH cartography. Accurate spatial reconstruction requires fitting to many spherical harmonics and is demonstrated using highly accurate Cauchy-characteristic numerical relativity waveforms. The loudest QNMs are mapped, and their reconstructed shapes are found to match the spheroidal harmonic predictions. The cartographic procedure is also applied to the quadratic QNMs – nonlinear features in the ringdown – and their reconstructed shapes are compared with expectations based on second-order perturbation theory. BH cartography allows us to determine the viewing angles that maximize the amplitude of the quadratic QNMs, an important guide for future searches, and is expected to lead to an improved understanding of nonlinearities in BH ringdown.
@@ I Introduction
The ground-based gravitational wave (GW) observatories LIGO [1] and Virgo [2] have already observed the merger of dozens of high-mass binary black holes (BHs) [3] and the current observing run is expected to yield many more observations [4, 5]. The final stage of the merger process, known as the ringdown, is associated with the remnant object settling down into its final state. The ringdown can be described by perturbation theory (PT), which predicts the existence of damped sinusoidal oscillations, known as quasinormal modes (QNMs) [6].
QNMs can be used directly for parameter estimation, especially for high-mass systems where only the late inspiral is observed [7, 8], or indirectly when incorporated as part of inspiral-merger-ringdown waveform models [9, 10, 11]. However, the most important application of QNMs is in tests of general relativity (GR), fundamental physics, and the nature of extremely compact objects. QNMs have been used to test Hawking’s area theorem and the consistency of the inspiral and ringdown signals [12, 13]. The no-hair theorem states that BHs should be completely characterized by their mass and spin (charge is excluded for astrophysical BHs). When multiple QNMs are detected, consistency between their frequencies allows us to test this in a program called BH spectroscopy [14, 15, 16]. This requires the clear identification of at least two QNMs and this has been the subject of much discussion in the literature [17, 18, 19, 20, 21]. However, to date, all observations are consistent with the standard picture of a vacuum Kerr BH in GR [22].
@@
Figure 1: Left: From NR, the spatial reconstruction of the fundamental QNM ($\ell=m=2$, $n=0$, prograde), normalized to agree with the corresponding spheroidal harmonic, plotted on the sphere in Mollweide projection. This mapping was performed for the 0001 Cauchy-characteristic extraction (CCE) waveform with a ringdown starting $t_{0}=8M$ after the peak of the $\ell=m=2$ mode strain (see Table 1). Right: From PT, the predicted spheroidal harmonic shape. In both panels, the real parts of the functions are plotted and they match well up to a rotation around the $z$-axis. The spatial mismatch between these two functions is $\mathcal{M}_{\alpha}=1.16\times 10^{-9}$ (see Sec. II.3). If the reconstruction is performed at different $t_{0}$, then the reconstruction is observed to rotate around the sphere with angular frequency equal to the real part of the QNM frequency. Because 0001 is nonprecessing, there is a symmetry between the ±mplus-or-$\pm m$ modes, and the $(\ell,m,n)=(2,-2,0)$ prograde QNM has a similar structure concentrated in the southern hemisphere.
@@
Another motivation for studying QQNMs is their possible relationship to the GW memory effect. Sometimes called non-linear displacement memory, this is the permanent displacement in test particle separation after the passage of a transient GW signal [33, 34, 35, 36, 37]. GW memory has not yet been observed, but is expected to be prominent in the azimuthally symmetric $m=0$ modes of the signal [38, 39]. Pairs of prograde modes with ±mplus-or-$\pm m$ indices have frequencies that are symmetrically placed in the complex plane and hence source a QQNM with a purely imaginary frequency. These nonoscillatory modes resemble a low-frequency, persistent displacement. Moreover, the combination of parent modes with ±mplus-or-$\pm m$ indices gives rise to QQNMs that are azimuthally symmetric with $m=0$. Together with the inherent nonlinear nature of QQNMs, this hints at a connection that might allow the GW memory effect to be studied using QNMs.
@@
To date, most fits (using either QNMs or QQNMs) have been performed to individual spherical harmonics of the waveform. Alternatively, the GW signal from a particular viewing direction can be fit. These approaches do not use all of the information in the NR simulation. Recently multimode, or angle-averaged, fits have been used, fitting a QNM model to all spherical modes simultaneously [40]. By using all the modes (effectively using the signal at all viewing directions) this exploits all of the information available in the NR simulation. Angle-averaged fits required another ingredient from PT: the predicted distribution of the GW radiation in different directions from the source. This distribution will be called the ‘spatial shape’ of the QNM and is a function on the 2-sphere. For the Kerr metric, the QNM shapes are given by the spin-weighted spheroidal harmonics [41]. This information is usually provided in the mode-mixing coefficients that relate the spheroidal and spherical harmonics.
@@
It is also possible to extract the QNM shapes directly from the NR data, without knowledge of the spheroidal harmonics. By fitting a feature with common time dependence to all spherical modes and treating the mode-mixing coefficients as a free parameters, the shape can be reconstructed numerically. In analogy to the identification of modes using their time dependence in BH spectroscopy, this approach is called BH cartography.
@@
This process can produce a spatial reconstruction of the shape of any QNM. An example of such a reconstruction is shown in Fig. 1 for the loudest fundamental QNM. Because this reconstruction is performed using the GW signal, it is formally defined on a sphere at future null infinity. However, it is expected that this will provide insight into the structure of the perturbations that source the GW signal and are generally considered to be located deep in the strong gravitational field, near the photon sphere of the remnant BH.
@@
BH cartography is a novel method for investigating the ringdown. Applied to the linear QNMs, it merely allows us to verify a well-understood prediction of first-order PT, namely the spheroidal harmonics. However, second-order perturbation theory, although capable of making unambiguous predictions for the metric perturbation, does not make such a clear and unique prediction for the QQNM shapes. There is no simple analog of the spheroidal harmonics at second order. This situation has led to different guesses, or predictions, being used for the QQNM mode shapes in the literature [30, 42]. Reconstructing the shapes of the QQNMs numerically allows us to test the different predictions. These reconstructions also allow us to determine the viewing directions that maximize the amplitude of the QQNMs with implications for their observability. Hence, knowing the quadratic mixing will be important for future waveform modeling.
@@
The layout of this paper is as follows. Sec. II describes the methods used, including a brief introduction to the key elements from BH PT (Sec. II.1), the least-squares fitting procedures to the NR data (Sec. II.2), the new techniques for spatially mapping components of the ringdown signal (Sec. II.3), and the Cauchy-characteristic extraction (CCE) waveforms used in this study (Sec. II.4). In Sec. III the mapping methods are demonstrated by applying them to the linear QNMs, which are already well understood from linear PT. This allows us to validate the BH cartography method. Then, Sec. IV applies the mapping to the QQNMs, revealing the spatial structure of the nonlinear perturbations. Concluding remarks are presented in Sec. V. Throughout, natural units are used in which $G=c=1$.
@@ II.1 Ingredients from perturbation theory
GW signals contain two polarizations, and these can be conveniently combined into a single complex quantity called simply the GW strain, $h=h_{+}-ih_{\times}$. The radially scaled GW strain can be expanded in spin-weight $s=-2$ spherical harmonics, ${}_{-2}Y^{\ell m}$,
@@
$rh^{(1)}$ $(t,\theta,\phi)=M\sum_{\ell,m}\sum_{n=0}^{\infty}$ (2)
[…]
$+$ $C^{\prime}_{\ell mn}e^{-i\omega^{\prime}_{\ell mn}(t-t_{0})}{}_{-2}S^{\prime}_{\ell m}(i\chi\omega^{\prime}_{\ell mn},\theta,\phi)\Big{]},$
where it has been assumed, for simplicity, that the frame $(\theta,\phi)$ is adapted to the remnant BH. That is, the spin of the remnant BH points along the $z$-axis of the chosen coordinate system. It is always possible to perform a rotation into such a frame. The superscript (1) in the notation is intended as a reminder that this part of the model for $h$ comes from linear PT. The ringdown is assumed to be the perturbation of a Kerr BH, with dimensionless spin $\chi$, which is naturally described using the spheroidal harmonics, ${}_{-2}S_{\ell m}(i\chi\omega_{\ell mn},\theta,\phi)$.
Each term in the expansion in Eq. 2 is a QNM. The indices $\ell\geq 2$ and $|m|\leq\ell$ label angular modes while the index $n\geq 0$ labels the radial-like overtones. For each triple $(\ell,m,n)$, there exists a pair of QNMs with different signs of the real part of their QNM frequency, $\omega^{\prime}_{\ell mn}=-\omega^{*}_{\ell-mn}$. The modes in this pair are sometimes referred to as the regular mode and mirror mode (denoted with a prime). The mirror modes are associated with the modified spheroidal harmonics, ${}_{-2}S^{\prime}_{\ell m}(i\chi\omega^{\prime}_{\ell mn},\theta,\phi)={}_{-2}S^{*}_{\ell-m}(i\chi\omega_{\ell-mn},\pi-\theta,\phi)$. In Eq. 2, this pair of regular and mirror modes has been written explicitly however, it will be convenient to distinguish them with another index $p\in\{+,-\}$ for the regular and mirror modes respectively. Therefore, a QNM is uniquely identified by the tuple of four indices $(\ell,m,n,p)$.
The distinction between regular and mirror modes is related to the direction in which the perturbation propagates around the Kerr BH. A mode is said to be prograde (i.e. co-rotating with the BH) if the signs of $m$ and $p$ are the same (if $mp>0$). If $mp<0$, then the mode is said to be retrograde (i.e. counter-rotating with the BH). Modes with $m=0$, along with all QNMs on Schwarzschild BHs, do not fit either the prograde or retrograde label. Typically, in a quasicircular merger, the prograde modes are excited with significantly larger amplitudes than the retrograde modes.
The QNM model in Eq. 2 involves many indices. In a moment, when QQNMs are introduced, the number of indices will increase further. Therefore, it is helpful to introduce the following abbreviated notation for the QNM model, and in particular for the spheroidal harmonics, in Eq. 2;
@@
The spheroidal harmonics are the natural basis for describing perturbations of Kerr BHs. The spheroidal basis does not align perfectly with the spherical basis. This leads to a phenomenon called mode mixing. The mode-mixing coefficients are functions of the BH spin and are defined as
@@
Going beyond linear PT introduces additional terms into the model in Eq. 3. Notable among these are the QQNMs. This second order part of the model, denoted with a superscript (2), is given by
@@ II.2 Least-squares fitting
The free parameters of the QNM model are the QNM complex amplitudes $C_{\alpha}$. The spheroidal harmonics and the QNM frequencies are functions of the final spin magnitude, $\chi_{f}$, of the remnant BH (the frequencies are also scaled by the remnant mass, $M_{f}$). The quantities $\chi_{f}$ and $M_{f}$ were obtained from the NR simulation data. Specifically, these were determined from the asymptotic Bondi data objects obtained using the scri package [44, 45, 46, 47] with the relevant files from the CCE data, and are given in Table 1.
To perform model fits to the data, we use the least-squares fit implemented in the qnmfits package [48]. More details on the single- and multi-mode fitting procedures are given in Appendix A. For all simulations, the time $t=0$ is set to be the peak of the amplitude of the 22 mode, $|\mathfrak{h}^{22}(t)|$.
For the NR data used in the fits, if the sums are truncated such that only modes with $\ell\leq\ell_{\rm max}$ are included, the resultant number of spherical harmonics ${}_{-2}Y^{\ell m}$ is $(\ell_{\rm max}+3)(\ell_{\rm max}-1)$. For the QNM models, if the sums are truncated such that all modes with $\ell\leq\ell_{\rm max}$ and $n\leq n_{\rm max}$ are included, the number of QNMs (including both $p=\pm 1$) is $2(\ell_{\rm max}+3)(\ell_{\rm max}-1)(n_{\rm max}+1)$.
Unless otherwise stated, all of the results in this paper were obtained using $\ell_{\rm max}=8$ and $n_{\rm max}=7$. Our method requires these maximum values to be large, but the precise values are unimportant. For overtones with $n\geq 8$, it becomes necessary to account for the QNM multiplet connected to the presence of the algebraically special QNM (see, for example, [49]). The qnm package by default caches mode-mixing coefficients up to the value $\ell_{\rm max}=8$. These were the practical reasons for the specific maximum values chosen.
The modeling approach described above including all the QNMs (up to a given $\ell_{\rm max}$ and $n_{\rm max}$), regardless of whether they are expected to be significantly excited, could be described as a “kitchen sink” method, i.e. throwing everything into the model. The obvious benefit of this approach is that no important modes can be accidentally left out of the model. The potential downside is the possibility of overfitting by the large number of free parameters associated with the QNMs with small amplitudes. This is not a problem for the cartographic reconstructions performed here. This is evidenced in later results by the stability of amplitudes over a wide range of start times and by their insensitivity to the precise choices of $\ell_{\rm max}$ and $n_{\rm max}$ (see, for example, the figure and discussion in Appendix B).